# Optimizing Large Language Model Training Using FP4 Quantization

### v1

@@ Appendix B Supplementary Proof for Differentiable Quantization Estimator
We present the complementary proof procedure for the Differentiable Gradient Estimator (DGE) method under actual quantization with vector-wise scaling factors. In the GeMM operation $Y=AW$, where $A$ is the activation tensor with dimensions ($s\times c_{i}$, sequence length × input channels) and $W$ is the weight tensor with dimensions ($c_{i}\times c_{o}$, input channels × output channels), quantization is applied along distinct dimensions to adhere to the mathematical logic of matrix multiplication. The quantization function is defined as:
For the weight tensor with dimensions ($c_{i}\times c_{o}$), channel-wise quantization is performed as follows:
@@
By applying the chain rule and referring to Equations 11 to 13, the relationship between $\partial{L}/\partial{W_{q}}$ and the actual weight gradient $\partial{L}/\partial{W}$ is established. According to Equation 13, the gradient $\partial L/\partial W_{q}^{\text{scaled}}$ can be expressed in terms of $\partial{L}/\partial{W_{q}}$ using the scaling factor $sf$:
@@
Appendix C Analyzing Quantization Difficulty Through Tensor Distribution

### v2

@@ C.2 Proof of DGE with Vector-wise Scaling Factors
Here we present the complementary proof procedure for the Differentiable Gradient Estimator (DGE) method under actual quantization with vector-wise scaling factors. In the GeMM operation $Y=AW$, where $A$ is the activation tensor with dimensions ($s\times c_{i}$, sequence length × input channels) and $W$ is the weight tensor with dimensions ($c_{i}\times c_{o}$, input channels × output channels), quantization is applied along distinct dimensions to adhere to the mathematical logic of matrix multiplication. For the weight tensor with dimensions ($c_{i}\times c_{o}$), channel-wise quantization is performed as follows:
@@
By applying the chain rule and referring to Equations 14 to 16, the relationship between $\partial{L}/\partial{W_{q}}$ and the actual weight gradient $\partial{L}/\partial{W}$ is established. According to Equation 16, the gradient $\partial L/\partial W_{q}^{\text{scaled}}$ can be expressed in terms of $\partial{L}/\partial{W_{q}}$ using the scaling factor $sf$:
@@
C.3 Mathematical Soundness of DGE Clipping
As introduced in Section 3.1, the differentiable quantization scheme includes a DGE correction term whose derivative can become unbounded near the midpoint of the quantization interval. Specifically, the derivative $f^{\prime}(x)$ is clipped to a maximum value of 3.0 in practice to prevent instability during training. In this appendix, we provide a detailed mathematical justification for this clipping operation and explain how a smoothing technique can be used to ensure continuity and numerical stability.
The derivative of the DGE approximation is given by:
$f^{\prime}(x)=\frac{1}{k}\cdot\left|\frac{2x}{\delta}-1\right|^{\frac{1}{k}-1}$ (23)
as also stated in Equation 8. This expression arises from differentiating the power-based approximation of the quantization function centered at $x=\delta/2$, where $\delta$ is the quantization range, and $k>1$ is a hyperparameter that controls the steepness of the transition.
The critical issue lies in the exponent $\frac{1}{k}-1$, which is negative for all $k>1$. Negative exponents in the form $|x|^{\alpha}$ with $\alpha<0$ are known to diverge as $x\to 0$, due to the reciprocal relationship. For instance, when $k=3$, the exponent becomes $-\frac{2}{3}$, leading to:
$f^{\prime}(x)=\frac{1}{3}\cdot\left|\frac{2x}{\delta}-1\right|^{-\frac{2}{3}}=\frac{1}{3\cdot\left(\left|\frac{2x}{\delta}-1\right|^{2/3}\right)}$ (24)
which becomes singular as $x\to\delta/2$. In this limit, the term $\left|\frac{2x}{\delta}-1\right|\to 0$, causing the denominator to vanish and $f^{\prime}(x)\to\infty$. This unbounded growth can lead to gradient explosions during optimization, especially when many input values are near $\delta/2$.
To eliminate the singularity at $x=\delta/2$, we redefine the absolute value function in a differentiable and bounded manner. Instead of using the sharp absolute value $|x|$, we adopt a smooth approximation defined as:
$|x|\approx\sqrt{x^{2}+\epsilon^{2}}$ (25)
where $\epsilon$ is a small positive constant. Substituting this into the derivative expression yields a smoothed version:
$f^{\prime}_{\text{smooth}}(x)=\frac{1}{k}\cdot\left(\sqrt{\left(\frac{2x}{\delta}-1\right)^{2}+\epsilon^{2}}\right)^{\frac{1}{k}-1}$ (26)
This formulation ensures that the denominator never reaches zero, as $\sqrt{(\cdot)^{2}+\epsilon^{2}}\geq\epsilon>0$, and therefore $f^{\prime}_{\text{smooth}}(x)$ remains bounded for all $x\in\mathbb{R}$. In effect, the singularity is replaced with a smooth transition that asymptotically approximates the original function as $\epsilon\to 0$, while maintaining differentiability and numerical stability during training.
We observe from the smoothed derivative expression presented in Equation 26 that the DGE correction term becomes bounded due to the regularization introduced by the epsilon term. Specifically:
$\lim_{x\to\delta/2}f^{\prime}_{\text{smooth}}(x)$ $=\lim_{x\to\delta/2}\frac{1}{k}\cdot\left(\sqrt{\left(\frac{2x}{\delta}-1\right)^{2}+\epsilon^{2}}\right)^{\frac{1}{k}-1}$ (27)
$=\frac{1}{k}\cdot\left(\sqrt{0+\epsilon^{2}}\right)^{\frac{1}{k}-1}=\frac{1}{k}\cdot\epsilon^{\frac{1}{k}-1}$ (28)
In practice, the DGE correction item is clipped with a predefined value, which is proved in Equation 28 to be mathematically equivalent since $k$ and $\epsilon$ are all constents. This approach is aligned with the simplicity principle in algorithm design: it avoids the need to modify the functional form while still achieving equivalent mathematical behavior around the singular point. As demonstrated above, both the smoothed and clipped versions share the key property of bounding the gradient magnitude in the vicinity of $x=\delta/2$, thereby ensuring stability during backpropagation.
Appendix D Analyzing Quantization Difficulty Through Tensor Distribution